# Directional expansiveness for $\mathbb{R}^{d}$-actions and for Penrose tilings

### v1

@@ Abstract.
We define and study two kinds of directional expansiveness, weak and strong, for an action $T$ of $\mathbb{R}^{d}$ on a compact metric space $X$. We show that for $\mathbb{R}^{2}$ finite local complexity (FLC) tiling dynamical systems, weak and strong expansiveness are the same, and are both equivalent to a simple coding property. Then we show for the Penrose tiling dynamical system, which is FLC, there are exactly five non expansive directions, the directions perpendicular to the 5th roots of unity. We also study Raphael Robinson’s set of $24$ Penrose Wang tiles and show the corresponding Penrose Wang tile dynamical system is strictly ergodic. Finally, we study two deformations of the Penrose Wang tile system, one where the square Wang tiles are all deformed into a $52\pi/$ 5 rhombus, and another where they are deformed into a set of eleven tetragon tiles. We show both of these are topologically conjugate to the Penrose tiling dynamical system.
@@ 1. Introduction
Once the basic theory of directional expansiveness for $\mathbb{R}^{d}$-actions has been established, we move to our main study: describing the expansive directions for the Penrose tiling dynamical system Our main theorem is the following:
@@ 1.1. Conventions
We will identify a $d$-tuple ${{\boldsymbol{w}}}=(w_{0},\cdots,w_{d-1})\in\mathbb{R}^{d}$ with a $d\times 1$ column vector. The usual dot product is denoted ${\boldsymbol{w}}\cdot{\boldsymbol{v}}$. The vectors ${\boldsymbol{e}}_{0},{\boldsymbol{e}}_{1},\dots,{\boldsymbol{e}}_{d-1}$ denote the usual standard basis for $\mathbb{R}^{d}$. All norms are Euclidean, and for $d\geq 2$, $B_{r}({\boldsymbol{w}}):=\{{\boldsymbol{v}}\in\mathbb{R}^{d}:||{\boldsymbol{v}}-{\boldsymbol{w}}||<r\}$ denotes the open radius $r$ ball (or disc in the case $d=2$) in $\mathbb{R}^{d}$. For $d>1$, $\mathbb{T}^{d}:=\mathbb{R}^{d}/\mathbb{Z}^{d}$, with $\mathbb{T}:=\mathbb{R}/\mathbb{Z}$ and $S^{1}=\{z\in\mathbb{C}:|z|=1\}$. We often identify $\mathbb{T}^{d}$ with $[0,1)^{d}$. For ${\boldsymbol{w}}\in\mathbb{R}^{d}$ we write $\lfloor{\boldsymbol{w}}\rfloor\in\mathbb{Z}^{d}$ for the coordinate-wise floor, and define $\{{\boldsymbol{w}}\}:={\boldsymbol{w}}-\lfloor{\boldsymbol{w}}\rfloor\in[0,1)^{d}$. We let $\gamma=\frac{1+\sqrt{5}}{2}\sim 1.6180\dots$ and $\alpha=1/\gamma=\gamma-1\sim 0.6180\dots$.
@@ 1.2. Acknowledgements
Most of the results in this paper are from the first author’s 2021 Ph.D. dissertation [17] at the George Washington University. Theorem 1.1 answers a question Mike Boyle once asked the second author.
@@ 2.1. Group actions
A $G$-action $T$ on $X$ is a continuous mapping $T:G\times X\to X$, written $T^{\boldsymbol{g}}x:=T({\boldsymbol{g}},x)$, satisfying $T^{{\boldsymbol{g}}+{\boldsymbol{h}}}=T^{\boldsymbol{g}}\circ T^{\boldsymbol{h}}$ for all ${\boldsymbol{g}},{\boldsymbol{h}}\in G$. We assume the acting group $G$ is non-compact and think of a $G$-action as a dynamical system, denoted $(X,G,T)$ or sometimes $T$ for short. A $G$-action $T$ is free, if every $x\in X$ is aperiodic: $T^{\boldsymbol{g}}x=x$ implies ${\boldsymbol{g}}={{\boldsymbol{0}}}$, and locally free if there exists $\omega>0$ so that $T^{\boldsymbol{g}}x=x$ implies $||{\boldsymbol{g}}||>\omega$.
A dynamical system $(X,G,T)$ is minimal if any closed $T$-invariant subset $K\subseteq X$ satisfies either $K$ or $K^{c}$ is empty, or equivalently, there exists a dense orbit: $X=\overline{\{T^{\boldsymbol{g}}x:{\boldsymbol{g}}\in G\}}$, (see [7]). We say $(X,G,T)$ is uniquely ergodic if there is a unique $T$-invariant Borel probability measure $\mu$ on $X$. Strictly ergodic means both minimal and uniquely ergodic, either of which implies $T$ is free.
@@ 2.2. Kronecker dynamical systems
(See e.g., [30], [36]) The dual group $\widehat{G}$ of a locally compact metric abelian group $G$ (see e.g., [36]) is the locally compact metric abelian group of continuous homomorphisms ${\chi}:G\to S^{1}$. In particular, $\widehat{\mathbb{R}^{d}}=\mathbb{R}^{d}$ with $\chi_{\boldsymbol{h}}({\boldsymbol{t}})=e^{2\pi i({\boldsymbol{h}}\cdot{\boldsymbol{t}})}$, and $\widehat{\mathbb{Z}^{d}}=\mathbb{T}^{d}$ with $\chi_{\boldsymbol{h}}({\boldsymbol{n}})=e^{2\pi i({\boldsymbol{h}}\cdot{\boldsymbol{n}})}$, both uncountable. But the compact $\mathbb{T}^{d}$ has a countable discrete dual $\widehat{\mathbb{T}^{d}}=\mathbb{Z}^{d}$ with $\chi_{\boldsymbol{n}}({\boldsymbol{h}})=e^{2\pi i({\boldsymbol{h}}\cdot{\boldsymbol{n}})}$.
@@
For an infinite compact metric abelian group $Y$, let $i:G\to Y$ be a continuous injective homomorphism. We define an algebraic Kronecker dynamical system $(Y,G,K)$ to be the free $G$-action $K^{\boldsymbol{g}}y:=y+i({\boldsymbol{g}})$. It satisfies $\widehat{K}\cong\Sigma_{K}$ via $\chi\mapsto\chi\circ i$. Minimality is equivalent to $\overline{i(G)}=K$, or $\chi\mapsto\chi\circ i$ injective, (see [30]). More generally, we call $(X,G,T)$ Kronecker if it is topologically conjugate to an algebraic Kronecker dynamical system. For a free Kronecker system, the Halmos-von Neumann Theorem, (see e.g., [13]) says $\Sigma_{K}$ is a complete topological conjugacy invariant. Also, every countable dense subgroup $\widehat{Y}$ occurs as $\Sigma_{K}$ for some $K$.
@@ 2.3. $\mathbb{Z}^{2}$ subshifts
For finite alphabet ${\mathcal{A}}=\{0,1,\dots,n-1\}$, $n\geq 2$, we call the product topology compact set ${\mathcal{A}}^{\mathbb{Z}^{2}}=\{z=(z_{\boldsymbol{m}})_{{\boldsymbol{m}}\in\mathbb{Z}^{2}}:z_{\boldsymbol{m}}\in{\mathcal{A}}\}$ the $\mathbb{Z}^{2}$ full shift, equipped with the $\mathbb{Z}^{2}$ shift action $S$ defined $(S^{\boldsymbol{n}}z)_{\boldsymbol{m}}=z_{{\boldsymbol{m}}-{\boldsymbol{n}}}$. The $\mathbb{Z}^{2}$ full shift dynamical system is $({\mathcal{A}}^{\mathbb{Z}^{2}},\mathbb{Z}^{2},S)$, and a $\mathbb{Z}^{2}$ subshift is $(Z,\mathbb{Z}^{2},S)$ where $Z\subseteq{\mathcal{A}}^{\mathbb{Z}^{2}}$ is any closed and $S$-invariant.
Given a rectangle $B\subseteq\mathbb{Z}^{2}$, #$\#(B)>1$, a word on $B$ is $b=(b_{\boldsymbol{m}})_{{\boldsymbol{m}}\in B}\in{\mathcal{A}}^{B}$. We say the word $b=(b_{\boldsymbol{m}})_{{\boldsymbol{m}}\in B}$ occurs in $z\in{\mathcal{A}}^{\mathbb{Z}^{2}}$, denoted $b|z$, if $(z_{{\boldsymbol{m}}-{\boldsymbol{n}}_{0}})_{{\boldsymbol{m}}\in B}=(b_{\boldsymbol{m}})_{{\boldsymbol{m}}\in B}$ for some ${\boldsymbol{n}}_{0}\in\mathbb{Z}^{2}$. We call a subshift $Z$ a subshift of finite type (SFT) if there is a finite set ${\mathcal{F}}$ of forbidden words, so that $Z=\{z\in{\mathcal{A}}^{\mathbb{Z}^{2}}:b|z\text{\ implies\ }b\notin{\mathcal{F}}\}$. For a special case, called a $1$-step SFT, we let ${\mathcal{F}}={\mathcal{F}}_{0}\cup{\mathcal{F}}_{1}$ where ${\mathcal{F}}_{j}\subseteq{\mathcal{A}}^{\{{{\boldsymbol{0}}},{\boldsymbol{e}}_{j}\}}$. A $1$-step SFT can also be described in terms of permitted words ${\mathcal{P}}_{j}={\mathcal{A}}^{\{{{\boldsymbol{0}}},{\boldsymbol{e}}_{j}\}}\backslash{\mathcal{F}}_{j}$.
@@ 2.4. Expansiveness for $\mathbb{Z}^{d}$ actions
    There are also versions of (i ) and (ii ) where the conclusion is simply that $y=T^{{\boldsymbol{s}}_{0}}x$, i.e., with no condition on $||{\boldsymbol{s}}_{0}||$. Gura [12] calls (i ) “separating” for $d=1$, and shows the horocycle flow satisfies (i ) but not (ii ) . For $d=1$, Katok and Hasselblatt [19] call this versionn of (ii ) for $d=1$ “expansiveness”. All (weak and strong) are equivalent in the FLC case.
@@ 2.5. Directional Expansiveness for $\mathbb{Z}^{d}$
Let $(X,\mathbb{Z}^{d},T)$ be a $\mathbb{Z}^{d}$ action and let ${\mathcal{V}}\subseteq\mathbb{R}^{d}$ be a $1\leq e<d$ dimensional direction. Then $T$ is expansive in the direction ${\mathcal{V}}$ if there exists A $\delta>0$ and an $r>0$ such that if $x,y\in X$ satisfy $\rho(T^{\boldsymbol{n}}x,T^{\boldsymbol{n}}y)\leq\delta$ for all ${\boldsymbol{n}}\in\mathbb{Z}^{d}\cap{\mathcal{V}}^{r}$, then $x=y$.
The sets of $e$-dimensional expansive and non-expansive directions are denoted in [3] by $\mathbb{E}_{d}(T)$ and $\mathbb{N}_{d}(T)$. In the paragraph following (2-1), Boyle and Lind [3] give the following useful characterization of directional expansiveness for a subshift.
@@
Let $(Z,\mathbb{Z}^{d},S)$ be a $\mathbb{Z}^{d}$ subshift, and let ${\mathcal{V}}\subseteq\mathbb{R}^{d}$ be an $1\leq e<d$ dimensional direction. For $z\in Z$, let $z[{\mathcal{V}}^{r}]:=(x_{\boldsymbol{m}})_{{\boldsymbol{m}}\in\mathbb{Z}^{2}\cap{\mathcal{V}}^{r}}$. Then ${\mathcal{V}}$ is an expansive direction if and only if there is an $r>0$ so that $z[{\mathcal{V}}^{r}]=z^{\prime}[{\mathcal{V}}^{r}]$ for $z,z^{\prime}\in Z$ implies $z=z^{\prime}$. We say ${\mathcal{V}}^{r}$ determines $z$.
Definition 2.6.
We say ${\mathcal{L}}\subseteq\mathbb{Z}^{2}$ thick if $\pi_{0}({\mathcal{L}})=\mathbb{Z}$ and $\pi_{1}({\mathcal{L}})=\mathbb{Z}$, where $\pi_{0},\pi_{1}:\mathbb{Z}^{2}\to\mathbb{Z}$ are the coordinate projections.
@@
Let $(Z\otimes Z,\mathbb{Z}^{2},S\otimes S)$ be the tensor square of a $\mathbb{Z}$-subshift, and let ${\mathcal{L}}\subseteq\mathbb{Z}^{2}$ be a thick set. Then ${\mathcal{L}}$ determines every $z\otimes z^{\prime}\in Z\otimes Z$.
@@
To find $z_{n_{0}}\otimes z_{n_{1}}$ for arbitrary $(n_{0},n_{1})\in\mathbb{Z}^{2}$ we use the thick property to find $(n_{0},m_{1}),(m_{0},n_{1})\in{\mathcal{L}}$. We use the first factor of $z_{n_{0}}\otimes z^{\prime}_{m_{1}}$, and the second factor of $z_{m_{0}}\otimes z^{\prime}_{n_{1}}$ to make $z_{n_{0}}\otimes z^{\prime}_{n_{1}}$. ∎
@@
Now suppose ${\mathcal{V}}$ is neither vertical nor horizontal. Define $Q_{(n_{0},n_{1})}:=[n_{0}-\frac{1}{2},n_{0}+\frac{1}{2})\times[n_{1}-\frac{1}{2},n_{1}+\frac{1}{2})$. Choose $r>0$ large enough that :=${\mathcal{L}}:=\bigcup_{\{{\boldsymbol{n}}\in\mathbb{Z}^{2}:Q_{\boldsymbol{n}}\cap{\mathcal{V}}\neq\emptyset\}}Q_{\boldsymbol{n}}\subseteq{\mathcal{V}}^{r}$. Clearly ${\mathcal{L}}$ is thick, so ${\mathcal{V}}^{r}$ contains a thick set. Lemma 2.7 shows ${\mathcal{V}}^{r}$ determines any $z\otimes z^{\prime}$, and the result again follows from Lemma 1. ∎

### v2

@@ Abstract.
We define and study two kinds of directional expansiveness, weak and strong, for an action $T$ of $\mathbb{R}^{d}$ on a compact metric space $X$. We show that for $\mathbb{R}^{2}$ finite local complexity (FLC) tiling dynamical systems, weak and strong expansiveness are the same, and are both equivalent to a simple coding property. Then we show for the Penrose tiling dynamical system, which is FLC, there are exactly five non expansive directions, the directions perpendicular to the 5th roots of unity. We also study Raphael Robinson’s set of $24$ Penrose Wang tiles and show the corresponding Penrose Wang tile dynamical system is strictly ergodic. Finally, we study two deformations of the Penrose Wang tile system, one where the square Wang tiles are all deformed into a $52\pi/$ 5 rhombus, and another where they are deformed into a set of eleven tetragon tiles. We show both of these are topologically conjugate to the Penrose tiliing dynamical system.
@@ 1. Introduction
Once the basic theory of directional expansiveness for $\mathbb{R}^{d}$-actions has been established, we move to our main study: describing the expansive directions for the Penrose tiling dynamical system. Our main theorem is the following:
@@ 1.1. Conventions
We will identify a $d$-tuple ${{\boldsymbol{w}}}=(w_{0},w_{1},\cdots,w_{d-1})\in\mathbb{R}^{d}$ with a $d\times 1$ column vector $[w_{0}w_{1}\cdots w_{d-1}]^{t}$. The usual dot product is denoted by ${\boldsymbol{w}}\cdot{\boldsymbol{v}}$. The vectors ${\boldsymbol{e}}_{0},{\boldsymbol{e}}_{1},\dots,{\boldsymbol{e}}_{d-1}$ denote the usual standard basis for $\mathbb{R}^{d}$. All norms are Euclidean, and for $d\geq 2$, $B_{r}({\boldsymbol{w}}):=\{{\boldsymbol{v}}\in\mathbb{R}^{d}:||{\boldsymbol{v}}-{\boldsymbol{w}}||<r\}$ denotes the open radius $r$ ball (or disc in the case $d=2$) in $\mathbb{R}^{d}$. For $d>1$, $\mathbb{T}^{d}:=\mathbb{R}^{d}/\mathbb{Z}^{d}$, with $\mathbb{T}:=\mathbb{R}/\mathbb{Z}$ and $S^{1}=\{z\in\mathbb{C}:|z|=1\}$. We often identify $\mathbb{T}^{d}$ with $[0,1)^{d}$. For ${\boldsymbol{w}}\in\mathbb{R}^{d}$ we write $\lfloor{\boldsymbol{w}}\rfloor\in\mathbb{Z}^{d}$ for the coordinate-wise floor, and define $\{{\boldsymbol{w}}\}:={\boldsymbol{w}}-\lfloor{\boldsymbol{w}}\rfloor\in[0,1)^{d}$. Throughout, we let $\gamma:=\frac{1+\sqrt{5}}{2}\approx 1.6180\dots$ and $\alpha:=1/\gamma=\gamma-1\approx 0.6180\dots$.
@@ 1.2. Acknowledgements
Most of the results in this paper are from the first author’s 2021 Ph.D. dissertation [17] at the George Washington University. Theorem 1.1 answers a question Mike Boyle once asked the second author. The authors thank Sébastien Labbé for pointing out that Figure 10 was mislabeled in an earlier version.
@@ 2.1. Group actions
A $G$-action $T$ on $X$ is a continuous mapping $T:G\times X\to X$, written $T^{\boldsymbol{g}}x:=T({\boldsymbol{g}},x)$, satisfying $T^{\boldsymbol{0}}={\rm id}$ and $T^{{\boldsymbol{g}}+{\boldsymbol{h}}}=T^{\boldsymbol{g}}\circ T^{\boldsymbol{h}}$ for all ${\boldsymbol{g}},{\boldsymbol{h}}\in G$. We assume the acting group $G$ is non-compact and think of a $G$-action as a dynamical system, denoted $(X,G,T)$ or sometimes $T$ for short. A $G$-action $T$ is free, if every $x\in X$ is aperiodic: $T^{\boldsymbol{g}}x=x$ implies ${\boldsymbol{g}}={{\boldsymbol{0}}}$, and locally free if there exists a $\omega>0$ so that $T^{\boldsymbol{g}}x=x$ implies $||{\boldsymbol{g}}||>\omega$.
A dynamical system $(X,G,T)$ is minimal if any closed $T$-invariant subset $K\subseteq X$ satisfies either $K$ or $K^{c}$ is empty, or equivalently, $X=\overline{O_{T}(x)}$ for some $x\in X$, where $O_{T}(x):=\{T^{\boldsymbol{g}}x:{\boldsymbol{g}}\in G\}$ is the orbit of $x$ under $T$, (see [7]). We say $(X,G,T)$ is uniquely ergodic if there is a unique $T$-invariant Borel probability measure $\mu$ on $X$. Strictly ergodic means both minimal and uniquely ergodic, either of which implies $T$ is free.
@@ 2.2. Kronecker dynamical systems
(See e.g., [30], [36]) The dual group $\widehat{G}$ of a locally compact metric abelian group $G$ (see e.g., [36]) is the locally compact metric abelian group of continuous homomorphisms ${\chi}:G\to S^{1}$ (in the compact-open topology). In particular, $\widehat{\mathbb{R}^{d}}=\mathbb{R}^{d}$ with $\chi_{\boldsymbol{h}}({\boldsymbol{t}})=e^{2\pi i({\boldsymbol{h}}\cdot{\boldsymbol{t}})}$, and $\widehat{\mathbb{Z}^{d}}=\mathbb{T}^{d}$ with $\chi_{\boldsymbol{h}}({\boldsymbol{n}})=e^{2\pi i({\boldsymbol{h}}\cdot{\boldsymbol{n}})}$, both uncountable. But the compact group $\mathbb{T}^{d}$ has a countable discrete dual $\widehat{\mathbb{T}^{d}}=\mathbb{Z}^{d}$ with $\chi_{\boldsymbol{n}}({\boldsymbol{h}})=e^{2\pi i({\boldsymbol{h}}\cdot{\boldsymbol{n}})}$.
@@
For an infinite compact metric abelian group $Y$, let $i:G\to Y$ be a continuous injective homomorphism. We define an algebraic Kronecker dynamical system $(Y,G,K)$ to be the free $G$-action $K^{\boldsymbol{g}}y:=y+i({\boldsymbol{g}})$. It satisfies $\widehat{K}\cong\Sigma_{K}$ via $\chi\mapsto\chi\circ i$. Minimality is equivalent to $\overline{i(G)}=K$, or $\chi\mapsto\chi\circ i$ injective, (see [30]). More generally, we call $(X,G,T)$ Kronecker if it is topologically conjugate to an algebraic Kronecker dynamical system. For a free Kronecker system, the Halmos-von Neumann Theorem, (see e.g., [13]) says $\Sigma_{K}$ is a complete topological conjugacy invariant. Also, every countable dense subgroup $\widehat{Y}\subseteq\widehat{G}$ occurs as $\Sigma_{K}$ for some $K$.
@@ 2.3. $\mathbb{Z}^{2}$ subshifts
For finite alphabet ${\mathcal{A}}=\{0,1,\dots,n-1\}$, $n\geq 2$, we call the product topology compact set ${\mathcal{A}}^{\mathbb{Z}^{2}}=\{z=(z_{\boldsymbol{m}})_{{\boldsymbol{m}}\in\mathbb{Z}^{2}}:z_{\boldsymbol{m}}\in{\mathcal{A}}\}$ the $\mathbb{Z}^{2}$ full shift space, equipped with the $\mathbb{Z}^{2}$ shift action $S$ defined by $(S^{\boldsymbol{n}}z)_{\boldsymbol{m}}=z_{{\boldsymbol{m}}-{\boldsymbol{n}}}$. The $\mathbb{Z}^{2}$ full shift dynamical system is $({\mathcal{A}}^{\mathbb{Z}^{2}},\mathbb{Z}^{2},S)$, and a $\mathbb{Z}^{2}$ subshift is $(Z,\mathbb{Z}^{2},S)$ where $Z\subseteq{\mathcal{A}}^{\mathbb{Z}^{2}}$ is any closed and $S$-invariant.
Given a rectangle $B\subseteq\mathbb{Z}^{2}$, #$\#(B)>1$, a word on $B$ is $b=(b_{\boldsymbol{m}})_{{\boldsymbol{m}}\in B}\in{\mathcal{A}}^{B}$. We say the word $b=(b_{\boldsymbol{m}})_{{\boldsymbol{m}}\in B}$ occurs in $z\in{\mathcal{A}}^{\mathbb{Z}^{2}}$ if $(z_{{\boldsymbol{m}}-{\boldsymbol{n}}_{0}})_{{\boldsymbol{m}}\in B}=(b_{\boldsymbol{m}})_{{\boldsymbol{m}}\in B}$ for some ${\boldsymbol{n}}_{0}\in\mathbb{Z}^{2}$. We call a subshift $Z$ a subshift of finite type (SFT) if there is a finite set ${\mathcal{F}}$ of forbidden words, so that now word in ${\mathcal{F}}$ occurs in any $z\in Z$. For a special case, called a $1$-step SFT, we let ${\mathcal{F}}={\mathcal{F}}_{0}\cup{\mathcal{F}}_{1}$ where ${\mathcal{F}}_{j}\subseteq{\mathcal{A}}^{\{{{\boldsymbol{0}}},{\boldsymbol{e}}_{j}\}}$. A $1$-step SFT can also be described in terms of permitted words ${\mathcal{P}}_{j}={\mathcal{A}}^{\{{{\boldsymbol{0}}},{\boldsymbol{e}}_{j}\}}\backslash{\mathcal{F}}_{j}$.
@@ 2.4. Expansiveness for $\mathbb{Z}^{d}$ actions
    There are also versions of (i ) and (ii ) where the conclusion is simply that $y=T^{{\boldsymbol{s}}_{0}}x$, i.e., with no condition on $||{\boldsymbol{s}}_{0}||$. Gura [12] calls (i ) “separating” for $d=1$, and shows the horocycle flow satisfies (i ) but not (ii ) . For $d=1$, Katok and Hasselblatt [19] call this version of (ii ) for $d=1$ “expansiveness”. All (weak and strong) are equivalent in the FLC case.
@@ 2.5. Directional Expansiveness for $\mathbb{Z}^{d}$
Let $(X,\mathbb{Z}^{d},T)$ be a $\mathbb{Z}^{d}$ action and let ${\mathcal{V}}\subseteq\mathbb{R}^{d}$ be a $1\leq e<d$ dimensional direction. Then $T$ is expansive in the direction ${\mathcal{V}}$ if there exist a $\delta>0$ and an $r>0$ such that if $x,y\in X$ satisfy $\rho(T^{\boldsymbol{n}}x,T^{\boldsymbol{n}}y)\leq\delta$ for all ${\boldsymbol{n}}\in\mathbb{Z}^{d}\cap{\mathcal{V}}^{r}$, then $x=y$.
In [3], the sets of $e$-dimensional expansive and non-expansive directions are denoted by $\mathbb{E}_{d}(T)$ and $\mathbb{N}_{d}(T)$. Let $(Z,\mathbb{Z}^{d},S)$ be a $\mathbb{Z}^{d}$ subshift, and let $J\subseteq\mathbb{R}^{d}$. We say a set $J$ determines $Z$ if for $z,z^{\prime}\in Z$, $z[J]=z^{\prime}[J]$ implies $z=z^{\prime}$. In the paragraph following (2–1) in [3], Boyle and Lind give the following useful characterization of directional expansiveness for a subshift.
@@
For a $\mathbb{Z}^{d}$ subshift $(Z,\mathbb{Z}^{d},S)$ and an $e$-dimensional direction ${\mathcal{V}}\subseteq\mathbb{R}^{d}$, $1\leq e<d$, ${\mathcal{V}}$ is an expansive direction if and only if ${\mathcal{V}}^{r}$ determines $Z$ for some $r>0$.
@@
Let $(Z\otimes Z,\mathbb{Z}^{2},S\otimes S)$ be the tensor square of a $\mathbb{Z}$-subshift. For $J\subseteq\mathbb{Z}^{2}\subseteq\mathbb{R}^{2}$, suppose $\pi_{0}(J)=\mathbb{Z}$ and $\pi_{1}(J)=\mathbb{Z}$, where $\pi_{0},\pi_{1}:\mathbb{Z}^{2}\to\mathbb{Z}$ are the coordinate projections. Then $J$ determines $Z\otimes Z$.
@@
To find $z_{n_{0}}\otimes z^{\prime}_{n_{1}}$ for arbitrary $(n_{0},n_{1})\in\mathbb{Z}^{2}$ we use $\pi_{0}(J)=\mathbb{Z}$ and $\pi_{1}(J)=\mathbb{Z}$ to find $(n_{0},m_{1}),(m_{0},n_{1})\in J$. Then we use the first factor of $z_{n_{0}}\otimes z^{\prime}_{m_{1}}$, and the second factor of $z_{m_{0}}\otimes z^{\prime}_{n_{1}}$ to make $z_{n_{0}}\otimes z^{\prime}_{n_{1}}$. ∎
@@
Now suppose ${\mathcal{V}}$ is neither vertical nor horizontal. Define $Q_{(n_{0},n_{1})}:=[n_{0}-\frac{1}{2},n_{0}+\frac{1}{2})\times[n_{1}-\frac{1}{2},n_{1}+\frac{1}{2})$. Choose $r>0$ large enough that $J:=\bigcup_{\{{\boldsymbol{n}}\in\mathbb{Z}^{2}:Q_{\boldsymbol{n}}\cap{\mathcal{V}}\neq\emptyset\}}Q_{\boldsymbol{n}}\subseteq{\mathcal{V}}^{r}$. $\pi_{0}(J)=\mathbb{Z}$ and $\pi_{1}(J)=\mathbb{Z}$, so the same holds for ${\mathcal{V}}^{r}$. Lemma 2.6 shows ${\mathcal{V}}^{r}$ determines $Z\otimes Z$, and the result follows from Lemma 2.5. ∎